OpenStudy (precal):

Calculus Problem-Optimization Problem A rectangle is inscribe in a circle of radius 2m. Find the maximum area of this rectangle.

OpenStudy (precal):

I need to draw it

OpenStudy (anonymous):

Yes. I wish the drawing tool worked.....

OpenStudy (precal):

ok well basically it is a rectangle made up of 2 rectangles. The dimensions for each rectangle is x and y

OpenStudy (anonymous):

I have to go eat lunch, but the answer should come out to be a square. That's what generally has the largest surface area in these optimization problems.

OpenStudy (precal):

but I need to give the set up. I believe the solution is 10 x 10 if your advice works out since that is the only logical solution on my puzzle worksheet

OpenStudy (anonymous):

Can you post the diagram here?

OpenStudy (precal):

Not sure, I would have to scan it in.

OpenStudy (anonymous):

Remember that the diagonal of the rectangle must be less than or equal to the diameter of the circle.

OpenStudy (anonymous):

not sure this is what your on about, but its a square inside a circle? draw a cross with 90 degree angles between each line, and then join the ends of each line (where the radius' meet the circumference and you get the biggest square possible

OpenStudy (anonymous):

http://www.twiddla.com/1471081 have fun

OpenStudy (precal):

@Agent48 If I draw it there, can I post it here?

OpenStudy (precal):

http://www.twiddla.com/1471081

OpenStudy (anonymous):

I would say L^2 + W^2 = 4^2 A = L * W A = L * sqrt(4^2 - L^2) Then you maximize A based on L.

OpenStudy (precal):

Ok that is the best I can do to draw it. My primary equation is A=x^2y I am having problems getting my secondary equation

OpenStudy (anonymous):

Why would it be x^2y?

OpenStudy (precal):

because the maximum area is the rectangle and that is the primary equation

OpenStudy (precal):

A=2x(y) I forgot that is A=lw

OpenStudy (anonymous):

Okay, so you are making x be half the width?

OpenStudy (precal):

I am given a rectangle that is made up of 2 rectangles. The dimension of each rectangle is x and y. I wish the drawing tool was working

OpenStudy (precal):

Length is x + x and width is y

OpenStudy (anonymous):

Ok

OpenStudy (precal):

I am stuck on the secondary equation

OpenStudy (precal):

I thought the secondary should have to do with the circle since the rectangle is inscribed in the circle of radius 2

OpenStudy (anonymous):

It would be the fact that the diagonal of the whole rectangle less than or equal to the diameter or the circle.

OpenStudy (precal):

Seems the equation tool is not working as well

OpenStudy (anonymous):

Latex wouldn't show up anyway.

OpenStudy (precal):

A=4x time the square root of (4-x^2)

OpenStudy (anonymous):

How did you get that?

OpenStudy (precal):

From the puzzle worksheet, it is the only solution that works (ie letter works)

OpenStudy (anonymous):

What were your first two equations? A = 2x y and?

OpenStudy (precal):

Primary is A=(2x)y but I need a secondary to get rid of x or y

OpenStudy (precal):

I need to go scan these pictures and repost. Let me go do that. I will repost this.